Math 128, Modern Geometry

D. Joyce, Clark University

7 Sep 2005

Read for Friday. Begin chapter 4.

Due Friday. Exercises from chapter 3: exercises 1, 2, 5b, 6, 8, 10, 15.

Today.

Continue our investigation on the assumptions behind the Pythagorean theorem.

Look at wallpaper groups at

   http://www.clarku.edu/∼djoyce/wallpaper/

especially the pages on “transformations of the plane” and “transformation and symmetry groups.”

Study the stereographic projection. There’s a discussion in my website on Compass Geometry at

   http://aleph0.clarku.edu/∼djoyce/java/compass/

We’ll look at the pages on “summary of inversion” and “stereographic projection.”